Chapter 10 Homework Saved Citywide Company issues bonds with a par value of $83,000 on their stated issue date. The bonds matReq 1 to 3 Reg 4 Reg 5 Compute the price of the bonds as of their issue date. (Round intermediate calculations to the nearestReq 1 to 3 Reg 4 Reg 5 Prepare the journal entry to record the bonds issuance. (Round intermediate calculations to the neareTABLE B.1 Present Value of 1 p=1/(1+i) Rate Periods 1% 2% 3% 4% 5% 6% 7% 8% 9% 10% 12% 15% 0 0 0.9901 0.9803 0.9706 0.9610TABLE B.3 Present Value of an Annuity of 1 Rate Perlods 1% 2% 3% 4% 5% 6% 7% 8% 9% 10% 12% 15% No von WN 0.9901 1.9704 2.9410f=[(1 + i) – 19i TABLE B.45 Future Value of an Annuity of 1 Rate Periods 1% 2% 3% 4% 5% 6% 7% 8% 9% 10% 12% 15% O VOWN 1.000
Answer #1
Par(maturity)value Semiannual Semiaanual
rate cash payment
83,000 * 5.5% = 4565
Number of payments 18
whether the bonds are issued at par ,at a Premium
discount ,or at a premium?
Table values are Based on
n= 18
i= 5%
Cash flow table value Amount present value
Par(maturity)value 0.4155 * 83,000 = 34487
interest (annuity) 11.6896 * 4565 = 53363
price of bonds 87850
TR General journal Debit Credit
Cash 87850
premium on bonds payable 4850
bonds payable 83,000